OpenStudy (anonymous):

A 4.0 Kg block is being dragged by a student who finds that he must exert a force of 14.0 N to keep the block moving at a constant velocity. Calculate the coefficient of kinetic friction for the block and the floor.

OpenStudy (anonymous):

PLEASE HELP!

OpenStudy (snowfire):

Okay so do you know the normal force here?

OpenStudy (snowfire):

The normal force is the force that the floor needs to exert to keep the block from falling through the ground. Since the block is not moving up or down, it must be equal to the weight of the block, which is the force due to gravity downwards. What is the weight of the block?

OpenStudy (anonymous):

so do you add the wright of the block to the force of gravity?

OpenStudy (snowfire):

No, the weight is equal to the acceleration due to gravity, which is 9.81 m/s^2 times the mass in kg. I'm sure you would be okay just using 10 m/s^2 to make it simpler though.

OpenStudy (cybershadow):

Given F = 14 N Normal force wud be N = mg = 4 * 10 = 40 N and frictional force is f = uN (where u = coefficient of friction) also F-f = ma as it is moving with constant velocity therefore a = 0 so F - f = 0 F = f 14 = uN u = 14/N u = 14/40 u = 0.35 That how i think it shud be done,

OpenStudy (snowfire):

Yep that seems sound. I was hoping Anon would reply back so I could walk him through it, but that's a good explanation